# SigMA: Path Signatures and Multi-head Attention for Learning Parameters in fBm-driven SDEs

### v1

@@ Abstract
Keywords: fractional Brownian motion; Stochastic differential equations; non-Markovian processes; parameter estimation; Hurst parameter; path signatures; self-attention; deep learning.
@@ 1 Introduction
Stochastic differential equations (SDEs) driven by fractional Brownian motion (fBm) arise in domains characterized by long-range dependence as well as pathwise roughness. Such applications include rough volatility modeling in quantitative finance [gatheral2018volatility] and degradation dynamics in engineering systems such as Li-ion batteries [shao2021degradation, boros2024deep]. A central quantity in models driven by fBm is the Hurst parameter $H\in(0,1)$, which governs temporal memory, self-similarity, and the local regularity of sample paths. Accurate estimation of this parameter from observed time series is critical, but challenging due to the non-Markovian nature of fBm-driven models. Moreover, in many applications, multiple parameters must be estimated jointly, such as the Hurst exponent, volatility of volatility, correlation, and/or mean-reversion parameters in rough volatility models. This is especially relevant in settings where pricing and risk sensitivities depend on combinations of these parameters [bayer2016pricing, el2019characteristic].
For $H\neq\tfrac{1}{2}$, fBm-driven processes are inherently non-Markovian and do not admit a semimartingale structure. As a result, many classical parameter estimation techniques become inapplicable or computationally intractable beyond very specific settings. These challenges motivate the development of data-driven architectures that can extract path-level features robustly across a wide range of parameter regimes. Recent machine learning techniques, including deep feedforward neural networks (FNNs) [mukherjee2023hurst], convolutional neural networks (CNNs) [stone2020calibrating], and long short-term memory (LSTM) [boros2024deep], have been proposed to address these limitations. However, these approaches often lack robustness across different path regimes or do not exploit the intrinsic geometric structure of long-memory paths.
Path signatures, arising from rough path theory [lyons2014rough], offer a principled way to represent continuous-time trajectories in a coordinate-free manner. Signature features have been shown to form an expressive and universal representation for path-dependent functionals, and have been used successfully in classification, forecasting, and medical time series [Lyons_2019_DeepSignature, morrill2021neural]. Recent efforts have also combined signatures with transformer architectures for time series modeling [moreno2024rough]. However, existing approaches typically focus on prediction or classification, rather than on parameter inference in fractional stochastic models.
@@
To this end, we propose SigMA (Signature Multi-head Attention), a neural architecture that integrates path signature features with multi-head self-attention, enhanced by convolutional preprocessing and multilayer perceptron (MLP) layers. Unlike prior methods such as Deep Signature Networks [Lyons_2019_DeepSignature], LSTM-based estimators [boros2024deep], or vanilla Transformers [vaswani_2017], SigMA is specifically designed for parameter estimation in fBm-driven SDEs, across both rough and long-memory regimes.
@@
  • Quantifying long-range dependence in Li-ion battery degradation data [boros2024deep].
Through extensive numerical and statistical evaluation, we demonstrate that SigMA outperforms existing benchmarks—CNNs [stone2020calibrating], LSTMs [boros2024deep], deep signature models [Lyons_2019_DeepSignature], and vanilla Transformers [vaswani_2017]—in terms of accuracy, robustness, and model compactness. Our findings support that path signatures, when integrated with attention-based architectures, offer a robust and scalable solution to parameter estimation in stochastic systems with rough or persistent temporal behavior.
@@ 2.2.1 Fractional Brownian motion
For $H>\tfrac{1}{2}$, the increments exhibit long-range dependence; for $H<\tfrac{1}{2}$, they exhibit rough anti-persistence. Almost surely, the sample paths are Hölder continuous of any order $\gamma<H$ [biagini2008stochastic, nualart2006malliavin].
There are different integral representations of fBm, in particular the characterisation provided by [mandelbrot_1968_fbm] as
@@ 2.2.2 Fractional Ornstein–Uhlenbeck process
The fOU process [kleptsyna2002statistical, cheridito2003fractional] is the fractional analogue of the OU process, where the process, $\{X^{fOU}_{t}\}_{t\geq 0}$, solves
@@ 2.2.3 Rough Heston process
The rHeston model [el2019characteristic], a non-Markovian extension of the classical Heston model, is another popular rough stochastic volatility model. It is a mean-reverting process that effectively captures both the volatility skew and volatility smiles observed in financial markets. In this model, the variance process, $\{X_{t}^{\mathrm{rH}}\}_{t\geq 0}$, satisifies
@@ 2.3 Self-attention mechanism
Self-attention is a core component of Transformer architectures [vaswani_2017] and provides a mechanism for modeling dependencies between elements in a sequence, regardless of their relative positions. Formally, given an input sequence $\mathbf{X}^{(n)}=(X_{0},\ldots,X_{n-1})\in\mathbb{R}^{n\times d}$, the attention mechanism computes a new representation for each element based on a weighted aggregation of the entire sequence.
@@ 2.4 Path signatures
A path signature transform [Lyons_1998_decay] is a feature extraction technique for sequential data, particularly effective for multivariate time series analysis. It characterizes a path $X=\{x_{t}|x_{t}=(x_{t}^{1},\ldots,x_{t}^{d})\in\mathbb{R}^{d},\,t\in[a,\,b]\}$, $a,\,b\in\mathbb{R}$ through a sequence of iterated integrals, capturing complex temporal dependencies and interactions between different dimensions of the data. Formally, the signature in the Stratonovich sense is defined as follows
@@
   Proposition 2.1 (Uniqueness ([Lyons_2010_uniqueness])).
Let $\mathbf{X}^{(n)}={(X_{0},\,\ldots,\,X_{n-1})}^{T}\in\mathbb{R}^{n\times d}$ be a path stream, and define its time-augmented version as $\mathbf{\widehat{X}}^{(n)}:={((X_{0},\,t_{0}),\,\ldots,\,(X_{n-1},\,t_{n-1}))}^{T}\in\mathbb{R}^{n\times(d+1)}$. Then the signature $\text{Sig}(\mathbf{\widehat{X}}^{(n)})$ can uniquely determine $\mathbf{X}^{(n)}$ up to a translation.
Proposition 2.1 implies that there is no information lost in the signature transform. The time-augmented path is necessary because the signature transform is invariant to time reparameterizations, meaning it encodes only the data arrival order, not the precise timing.
   Proposition 2.2 (Factorial decay ([Lyons_1998_decay])).
@@
Proposition 2.2 indicates that the terms of the signature decay factorially in size. This ensures that the most significant contributions are given by the lower order terms which justifies the use of truncated signatures.
@@ 3.1 Convolutional feature extraction
Consider a data stream $\mathbf{X}^{(n)}={(X_{0},\,\ldots,\,X_{n-1})}^{T}\in\mathbb{R}^{n\times d}$, interpreted as a discretized path. It is beneficial to apply a feature mapping before computing the signature transform (see [Lyons_2019_DeepSignature]). Therefore, at the input of SigMA, a convolutional layer $\Phi(\mathbf{X}^{(n)};\,W^{\Phi},B^{\Phi}):\mathbb{R}^{n\times d}\mapsto\mathbb{R}^{\widetilde{n}\times\widetilde{d}}$ with kernel size $k\in\mathbb{N}^{+}$ and stride $s\in\mathbb{N}^{+}$ is applied to perform the feature mapping
@@
where $W^{\Phi}\in\mathbb{R}^{\widetilde{d}\times k\times d}$ and $B^{\Phi}\in\mathbb{R}^{\widetilde{d}}$ are the weights and biases of this convolutional layer, respectively, while $\widetilde{n}\in\mathbb{N}^{+}$ and $\widetilde{d}\in\mathbb{N}^{+}$ represent the dimensions of outputs. We denote $\star$ as the valid cross-correlation operator and $\mathbf{1}_{\widetilde{n}}\in\mathbb{R}^{\widetilde{n}}$ is an all-ones column vector. This convolutional layer is shown in Section 4.2.3 to improve the accuracy of the SigMA model. After the feature mapping, according to Proposition 2.1, it is beneficial to augment the output with time before taking the signature transform.
@@ 3.2 Signature transform on lifted paths
Taking the signature transform can extract the geometric features from original paths while filtering out irrelevant information, however, it will consume the stream-like nature of the input data (see [Lyons_2019_DeepSignature]). Let us consider a stream of data $\mathbf{Y}^{(\widetilde{n})}={(Y_{0},\,\ldots,\,Y_{\widetilde{n}-1})}^{T}\in\mathbb{R}^{\widetilde{n}\times\widetilde{d}}$, which is the output of the convolutional layer and has been augmented. Here, $\widetilde{d}$ denotes the number of features, $\widetilde{n}$ denotes the length of the sequence and is assumed to be even.
@@ 3.3 Multi-head self-attention on signature features
Recall that the signature transform can be considered as a feature mapping which helps to extract the information from the input data while filtering out irrelevant information. The signature transform yields a high-dimensional feature representation that typically requires additional processing. A common approach is to use a FNN, as in Deep Signature Networks [Lyons_2019_DeepSignature]. In addition to an FNN, we incorporate a self-attention layer, which is particularly effective for joint multi-parameter estimation.
@@
where $\text{head}_{i}=\text{Self-Attention}({(\text{Sig}_{i}^{N}(\mathbf{Y}^{(\widetilde{n})}_{0:\frac{\widetilde{n}}{2}}),\,\text{Sig}_{i}^{N}(\mathbf{Y}^{(\widetilde{n})}_{0:\widetilde{n}}))}^{T})$ for $i=1,\,\ldots,\,N$. Here we also omit all the weight matrix except the linear projection matrix $\widetilde{W}_{O}$ for the output. After the $N$-head self-attention layer, we need to project the output to an appropriate dimension using another MLP architecture, which has been shown to be important for the robustness of the SigMA model in Section 4.2.3.
@@ 4 Numerical and empirical experiments
In this section, we conduct experiments to verify the advantages of the architecture described in Section 3 and motivate the architectural choices. To this end, we compare SigMA [^3] Source code is available on https://github.com/changanluoxue/SigMA.git with four other neural network-based models. The first model is a CNN proposed by [stone2020calibrating]. [^4] Source code is available on https://github.com/henrymstone/rough-calibration-using-CNNS The second model is a signature-based (DeepSigNet) model proposed by [Lyons_2019_DeepSignature]. [^5] Source code is available on https://github.com/patrick-kidger/Deep-Signature-Transforms The third model is a modified Transformer based on [vaswani_2017]. The fourth model is a LSTM model proposed by [csanady2024parameter]. [^6] Source code is available on https://github.com/aielte-research/LMSParEst The specific architectures of these four models are detailed in Appendix A.
This section is divided into five subsections. Firstly, we clarify the synthetic data generation and some implementation details. In Section 4.1, we perform sensitivity analysis to determine the optimal architecture of the SigMA model. Next, we discuss the single and multiple parameters estimation problems in Sections 4.3 and  4.4. Finally, in Section 4.5, we present two empirical studies: one using realized volatility data from financial markets and another employing Li-ion battery degradation data.
@@ 4.1 Data generation and implementation details
We use the Cholesky decomposition to simulate fBm and subsequently generate fOU paths with $\alpha=0.5$, $\mu=0.15$, $\sigma=0.2$, and $X_{0}^{fOU}=0.01$. For simulating rHeston paths, we apply the fast algorithm proposed by [ma_2022_fast], using the parameters $\kappa_{1}=0.1$, $\kappa_{2}=0.03$, $\theta=0.3$, and $X_{0}^{rH}=0.01$.
@@ 4.2.2 Effect of signature truncation order
In this section, we aim to explore the effects of varying signature truncation orders on the accuracy of the SigMA model. In this experiment, we use the simulated paths from Section 4.1 with lengths from $\{100,\,500\}$ and we vary the signature truncation order from $\{1,\,3,\,5\}$, while keeping other settings consistent with those in Section 4.1. In addition to the SigMA model, we introduce a simpler version with a single-head self-attention layer, called SigSA (details in Appendix A). We also present results for the DeepSigNet model because it also incorporates the signature transform. Based on the finding from the previous section, we set a large stride as half of the input length.
@@
(a) The input lengths equal to $100$
(b) The input lengths equal to $500$
Figure 2: Test RMSE trends for Hurst estimation across stochastic processes (Left: fBms; Middle: fOU; Right: rHeston) of varying lengths (100-500), obtained using the signature-based $N$ models with varying signature truncation orders (1-5), are averaged over 3 training runs. Here, $H\sim\text{Uniform}(0,\,1)$ for the fBm, $H\sim\text{Uniform}(0.5,\,1)$ for the fOU, and $H\sim\text{Beta}(1,\,9)$ for the rHeston.
Table 2: Test RMSEs for Hurst estimation across stochastic processes of varying lengths (100-500), obtained using the signature-based $N$ models with varying signature truncation orders (1-5), are averaged over 3 training runs. Here, $H\sim\text{Uniform}(0,\,1)$ for the fBm, $H\sim\text{Uniform}(0.5,\,1)$ for the fOU, and $H\sim\text{Beta}(1,\,9)$ for the rHeston. The best accuracies across different signature truncation orders are in bold.

### v2

@@ Abstract
keywords:
fractional Brownian motion , stochastic differential equations , non-Markovian processes , parameter estimation , path signatures , self-attention
journal: Neurocomputing
[…]
@@ 1 Introduction
Stochastic differential equations (SDEs) driven by fractional Brownian motion (fBm) arise in domains characterized by long-range dependence as well as pathwise roughness. Such applications include rough volatility modeling in quantitative finance [12] and degradation dynamics in engineering systems such as Li-ion batteries [29, 3]. A central quantity in models driven by fBm is the Hurst parameter $H\in(0,1)$, which governs temporal memory, self-similarity, and the local regularity of sample paths. Accurate estimation of this parameter from observed time series is critical, but challenging due to the non-Markovian nature of fBm-driven models. Moreover, in many applications, multiple parameters must be estimated jointly, such as the Hurst exponent, volatility of volatility, correlation, and/or mean-reversion parameters in rough volatility models. This is especially relevant in settings where pricing and risk sensitivities depend on combinations of these parameters [1, 10].
For $H\neq\tfrac{1}{2}$, fBm-driven processes are inherently non-Markovian and do not admit a semimartingale structure. As a result, many classical parameter estimation techniques become inapplicable or computationally intractable beyond very specific settings. These challenges motivate the development of data-driven architectures that can extract path-level features robustly across a wide range of parameter regimes. Recent machine learning techniques, including deep feedforward neural networks (FNNs) [26], convolutional neural networks (CNNs) [30], and long short-term memory (LSTM) [3], have been proposed to address these limitations. However, these approaches often lack robustness across different path regimes or do not exploit the intrinsic geometric structure of long-memory paths.
Path signatures, arising from rough path theory [20], offer a principled way to represent continuous-time trajectories in a coordinate-free manner. Signature features have been shown to form an expressive and universal representation for path-dependent functionals, and have been used successfully in classification, forecasting, and medical time series [15, 25]. Recent efforts have also combined signatures with transformer architectures for time series modeling [24]. However, existing approaches typically focus on prediction or classification, rather than on parameter inference in fractional stochastic models.
@@
To this end, we propose SigMA (Signature Multi-head Attention), a neural architecture that integrates path signature features with multi-head self-attention, enhanced by convolutional preprocessing and multilayer perceptron (MLP) layers. Unlike prior methods such as Deep Signature Networks [15], LSTM-based estimators [3], or vanilla Transformers [32], SigMA is specifically designed for parameter estimation in fBm-driven SDEs, across both rough and long-memory regimes.
@@
  • Quantifying long-range dependence in Li-ion battery degradation data [3].
Through extensive numerical and statistical evaluation, we demonstrate that SigMA outperforms existing benchmarks—CNNs [30], LSTMs [3], deep signature models [15], and vanilla Transformers [32]—in terms of accuracy, robustness, and model compactness. Our findings support that path signatures, when integrated with attention-based architectures, offer a robust and scalable solution to parameter estimation in stochastic systems with rough or persistent temporal behavior.
@@ 2.2.1 Fractional Brownian motion
For $H>\tfrac{1}{2}$, the increments exhibit long-range dependence; for $H<\tfrac{1}{2}$, they exhibit rough anti-persistence. Almost surely, the sample paths are Hölder continuous of any order $\gamma<H$ [2, 27].
There are different integral representations of fBm, in particular the characterisation provided by [22] as
@@ 2.2.2 Fractional Ornstein–Uhlenbeck process
The fOU process [17, 5] is the fractional analogue of the OU process, where the process, $\{X^{fOU}_{t}\}_{t\geq 0}$, solves
@@ 2.2.3 Rough Heston process
The rHeston model [10], a non-Markovian extension of the classical Heston model, is another popular rough stochastic volatility model. It is a mean-reverting process that effectively captures both the volatility skew and volatility smiles observed in financial markets. In this model, the variance process, $\{X_{t}^{\mathrm{rH}}\}_{t\geq 0}$, satisifies
@@ 2.3 Self-attention mechanism
Self-attention is a core component of Transformer architectures [32] and provides a mechanism for modeling dependencies between elements in a sequence, regardless of their relative positions. Formally, given an input sequence $\mathbf{X}^{(n)}=(X_{0},\ldots,X_{n-1})\in\mathbb{R}^{n\times d}$, the attention mechanism computes a new representation for each element based on a weighted aggregation of the entire sequence.
@@ 2.4 Path signatures
A path signature transform [19] is a feature extraction technique for sequential data, particularly effective for multivariate time series analysis. It characterizes a path $X=\{x_{t}|x_{t}=(x_{t}^{1},\ldots,x_{t}^{d})\in\mathbb{R}^{d},\,t\in[a,\,b]\}$, $a,\,b\in\mathbb{R}$ through a sequence of iterated integrals, capturing complex temporal dependencies and interactions between different dimensions of the data. Formally, the signature in the Stratonovich sense is defined as follows
@@
Proposition 1 (Uniqueness ([13]))
Let $\mathbf{X}^{(n)}={(X_{0},\,\ldots,\,X_{n-1})}^{T}\in\mathbb{R}^{n\times d}$ be a path stream, and define its time-augmented version as
$\mathbf{\widehat{X}}^{(n)}:={((X_{0},\,t_{0}),\,\ldots,\,(X_{n-1},\,t_{n-1}))}^{T}\in\mathbb{R}^{n\times(d+1)}.$
Then the signature $\text{Sig}(\mathbf{\widehat{X}}^{(n)})$ can uniquely determine $\mathbf{X}^{(n)}$ up to a translation.
Proposition 1 implies that there is no information lost in the signature transform. The time-augmented path is necessary because the signature transform is invariant to time reparameterizations, meaning it encodes only the data arrival order, not the precise timing.
Proposition 2 (Factorial decay ([19]))
@@
Proposition 2 indicates that the terms of the signature decay factorially in size. This ensures that the most significant contributions are given by the lower order terms which justifies the use of truncated signatures.
Proposition 3 (Chen’s identity ([4]))
Let $X:[s,t]\to\mathbb{R}^{d}$ be a path of bounded variation, and let $u\in(s,t)$ be an intermediate time. The signature of the entire path $\text{Sig}(X_{[s,t]})$ satisfies
$\text{Sig}(X_{[s,t]})=\text{Sig}(X_{[s,u]})\otimes\text{Sig}(X_{[u,t]}).$
Proposition 3 shows that the signature of a long trajectory implicitly contains the signatures of its subpaths. In practice, this algebraic decomposition of path signatures justifies the ”lifting” procedure within SigMA, as detailed in Section 3.2.
@@ 3.1 Convolutional feature extraction
Consider a data stream $\mathbf{X}^{(n)}={(X_{0},\,\ldots,\,X_{n-1})}^{T}\in\mathbb{R}^{n\times d}$, interpreted as a discretized path. It is beneficial to apply a feature mapping before computing the signature transform (see [15]). Therefore, at the input of SigMA, a convolutional layer $\Phi(\mathbf{X}^{(n)};\,W^{\Phi},B^{\Phi}):\mathbb{R}^{n\times d}\mapsto\mathbb{R}^{\widetilde{n}\times\widetilde{d}}$ with kernel size $k\in\mathbb{N}^{+}$ and stride $s\in\mathbb{N}^{+}$ is applied to perform the feature mapping
@@
where $W^{\Phi}\in\mathbb{R}^{\widetilde{d}\times k\times d}$ and $B^{\Phi}\in\mathbb{R}^{\widetilde{d}}$ are the weights and biases of this convolutional layer, respectively, while $\widetilde{n}\in\mathbb{N}^{+}$ and $\widetilde{d}\in\mathbb{N}^{+}$ represent the dimensions of outputs. We denote $\star$ as the valid cross-correlation operator and $\mathbf{1}_{\widetilde{n}}\in\mathbb{R}^{\widetilde{n}}$ is an all-ones column vector. This convolutional layer is shown in Section 4.2.3 to improve the accuracy of the SigMA model. After the feature mapping, according to Proposition 1, it is beneficial to augment the output with time before taking the signature transform. This augmentation enables the signature to encode the exact arrival times and the specific $\Delta t$ between observations. Consequently, SigMA can natively process irregularly sampled data and adapt to varying sampling frequencies.
Remark 1
The convolutional layer in SigMA serves not only as a feature encoder but also enables SigMA to handle high-dimensional processes. For a path of dimension $d$, the number of terms in its signature truncated at order $N$ grows exponentially as $O(d^{N})$. If the signature were computed directly on high-dimensional raw data, the computational cost would become expensive. SigMA mitigates this challenge by first projecting the raw path $\mathbf{X}^{(n)}\in\mathbb{R}^{n\times d}$ into a lower-dimensional latent representation of size $\tilde{d}$ before computing signatures. Therefore, the subsequent signature transform is independent of the raw input dimension, and the size of the truncated signature scales as $O(\tilde{d}^{N})$. This decouples the exponential growth of the signature tensor from the original input dimension, enabling SigMA to scale efficiently to higher-dimensional SDEs without causing a parameter explosion in downstream layers.
@@ 3.2 Signature transform on lifted paths
Taking the signature transform can extract the geometric features from original paths while filtering out irrelevant information, however, it will consume the stream-like nature of the input data (see [15]). Let us consider a stream of data $\mathbf{Y}^{(\widetilde{n})}={(Y_{0},\,\ldots,\,Y_{\widetilde{n}-1})}^{T}\in\mathbb{R}^{\widetilde{n}\times\widetilde{d}}$, which is the output of the convolutional layer and has been augmented. Here, $\widetilde{d}$ denotes the number of features, $\widetilde{n}$ denotes the length of the sequence and is assumed to be even.
@@
As shown in Equation (2.4), the path signature is defined as global iterated integration. However, when a truncated signature is computed over the entire trajectory, geometric information from different temporal regions becomes compressed into a single feature vector. By introducing lifted segments and computing signatures on these subsegments separately, the model receives explicit access to both local and global geometric features of the path. This provides a practical mechanism for constructing a multi-scale representation of the trajectory. Such segmented or streaming applications of signatures are commonly used in machine learning settings, where signatures are computed on successive windows of a time series in order to capture localized geometric information [6, 15].
@@ 3.3 Multi-head self-attention on signature features
Recall that the signature transform can be considered as a feature mapping which helps to extract the information from the input data while filtering out irrelevant information. The signature transform yields a high-dimensional feature representation that typically requires additional processing. A common approach is to use a FNN, as in Deep Signature Networks [15]. In addition to an FNN, we incorporate a self-attention layer, which is particularly effective for joint multi-parameter estimation.
@@
where $\text{head}_{i}=\mathcal{A}({(\text{Sig}_{i}^{N}(\mathbf{Y}^{(\widetilde{n})}_{0:\frac{\widetilde{n}}{2}}),\,\text{Sig}_{i}^{N}(\mathbf{Y}^{(\widetilde{n})}_{0:\widetilde{n}}))}^{T})$ for $i=1,\,\ldots,\,N$. Here we also omit all the weight matrix except the linear projection matrix $\widetilde{W}_{O}$ for the output.
The design of the attention mechanism is motivated by structural properties of the signature transform. Path signatures form a graded tensor algebra representation of trajectories, where different truncation levels correspond to geometric interactions of different orders [18, 11]. Functions of path signatures are known to be universal approximators for functionals of streams [16], and the expected signature uniquely characterizes the law of stochastic processes under mild conditions [7]. These results provide a theoretical justification for using signature features for parameter inference. In practice, truncated signatures produce a hierarchy of features with rapidly decaying magnitude across orders. Allowing the model to process these levels separately enables the architecture to assign different importance weights to different interaction orders. The multi-head attention layer therefore acts as an adaptive mechanism that learns which signature levels are most informative for the inference task. While this mechanism is primarily motivated by these structural properties, its effectiveness is ultimately validated empirically in the experiments.
After the $N$-head self-attention layer, we need to project the output to an appropriate dimension using another multilayer perceptron (MLP) architecture, which has been shown to be important for the robustness of the SigMA model in Section 4.2.3.
@@ 4 Numerical and empirical experiments
In this section, we conduct experiments to verify the advantages of the architecture described in Section 3 and motivate the architectural choices. To this end, we compare SigMA [^3] Source code is available on https://github.com/changanluoxue/SigMA.git with four other neural network-based models. The first model is a CNN proposed by [30]. [^4] Source code is available on https://github.com/henrymstone/rough-calibration-using-CNNS The second model is a signature-based (DeepSigNet) model proposed by [15]. [^5] Source code is available on https://github.com/patrick-kidger/Deep-Signature-Transforms The third model is a modified Transformer based on [32]. The fourth model is a LSTM model proposed by [8]. [^6] Source code is available on https://github.com/aielte-research/LMSParEst The specific architectures of these four models are detailed in A.
This section is divided into six subsections. Firstly, we clarify the synthetic data generation and some implementation details. In Section 4.1, we perform sensitivity analysis to determine the optimal architecture of the SigMA model. Next, we discuss the single and multiple parameters estimation problems in Sections 4.3 and  4.4. In Section 4.5, we present two empirical studies: one using realized volatility data from financial markets and another employing Li-ion battery degradation data. Finally, based on the presented results, we summarize the differences between SigMA and comparable models in terms of architectural designs, applicable scenarios, and performance in Section 4.6.
@@ 4.1 Data generation and implementation details
We report the number of trainable parameters for all neural architectures as the hardware-independent proxy for computational complexity and memory footprint. We use the Cholesky decomposition to simulate fBm and subsequently generate fOU paths with $\alpha=0.5$, $\mu=0.15$, $\sigma=0.2$, and $X_{0}^{fOU}=0.01$. For simulating rHeston paths, we apply the fast algorithm proposed by [21], using the parameters $\kappa_{1}=0.1$, $\kappa_{2}=0.03$, $\theta=0.3$, and $X_{0}^{rH}=0.01$.
@@ 4.2.2 Effect of signature truncation order
In this section, we aim to explore the effects of varying signature truncation orders on the accuracy of the SigMA model. In this experiment, we use the simulated paths from Section 4.1 with lengths from $\{100,\,500\}$ and we vary the signature truncation order from $\{1,\,3,\,5\}$, while keeping other settings consistent with those in Section 4.1. In addition to the SigMA model, we introduce a simpler version with a single-head self-attention layer, called SigSA (details in A). We also present results for the DeepSigNet model because it also incorporates the signature transform. Based on the finding from the previous section, we set a large stride as half of the input length.
@@
Table 2: Test RMSEs for Hurst estimation across stochastic processes of varying lengths (100-500), obtained using the signature-based NN models with varying signature truncation orders (1-5), are averaged over 3 training runs. Here, $H\sim\text{Uniform}(0,\,1)$ for the fBm, $H\sim\text{Uniform}(0.5,\,1)$ for the fOU, and $H\sim\text{Beta}(1,\,9)$ for the rHeston. The best accuracies across different signature truncation orders are in bold.
@@
(a) The input lengths equal to $100$
(b) The input lengths equal to $500$
Figure 2: Test RMSE trends for Hurst estimation across stochastic processes (Left: fBms; Middle: fOU; Right: rHeston) of varying lengths (100-500), obtained using the signature-based NN models with varying signature truncation orders (1-5), are averaged over 3 training runs. Here, $H\sim\text{Uniform}(0,\,1)$ for the fBm, $H\sim\text{Uniform}(0.5,\,1)$ for the fOU, and $H\sim\text{Beta}(1,\,9)$ for the rHeston.
For further explanations, we perform an interpretability analysis based on attention weights and MLP feature importance. Specifically, we introduce two metrics: a segment-wise index averages the softmax weights from query-key interactions to measure the contribution of different lifted segments, and a level-wise index uses the mean absolute weights of the MLP’s first linear layer to measure the importance of different signature truncation levels. The results, shown in Figure 3, suggest that both local and global segments contribute comparably to the prediction, while the most informative signature level appears around truncation order $N=3$. These observations are consistent with the approximate scale-invariant properties of the processes considered.
(a) Segment-wise index
(b) Level-wise index
Figure 3: Interpretability analysis for SigMA on rHeston paths with length 500.
One potential concern is the computational cost of evaluating path signatures, since their mathematical definition involves iterated integrals (see Equation (2.4)). However, thanks to Chen’s identity (Proposition 3), the computation can be performed algebraically in practice, avoiding the need for explicit numerical integration. In particular, the signature of a piecewise linear path can be computed by first evaluating the signatures of linear segments between consecutive observations and then combining them recursively through tensor products.
In our implementation, this procedure is carried out using the highly optimized iisignature package [28], which provides efficient algorithms for computing truncated signatures and log-signatures. For a fixed truncation order, the computational cost of signature evaluation scales approximately linearly with the sequence length, making it practical for the trajectory lengths considered in our experiments.